# Bridging the Gulf of Envisioning: Cognitive Design Challenges in LLM Interfaces

### v1

@@ 6. Discussion
This work argues that HCI designers must now tackle not just system mental models, but additional cognitive models to capture people’s understanding of cognitive tasks. Thanks to LLMs, people are now able to describe how they think about satisfying their task goals and ask systems to do it for them. At the same time, the specification of actions needed to implement these task processes in LLM systems has virtually disappeared from the design of interfaces. It is no longer necessary to translate intentions into action specifications in the language of a system; indeed, that is impossible because the LLM’s internal processing language is indiscernible. Further, users’ system mental models for LLMs are virtually non-existent. Even experts can’t look at the operating LLM and identify how its training, database, and operation produce specific mappings from input to output. The user will have no better luck in attempting to predict LLM output given a prompt they wish to try.
With the absence of these action-based elements from interaction design, we identify new opportunities for specifying actions through intentions. With the accumulated database of texts encoded into LLMs through training, an implicit map emerges of how past goals were developed through intentions and executed to produce acceptable output. To access the valuable information encoded in an LLM, users must envision the original creator’s goal and intentions; through their identification, the connections to related words in the LLM produce output with similar value. It is only within the context of intentions that people can understand the value of generation and evaluate its output. This mental simulation of how one might perform the generation task produces insight into the implicit qualities of past task performers and their successful product (by virtue of appearing in the corpus). As this changes and more AI-generated texts populate further training corpora, connections through intention will be less successful. However, the value of generated texts for users will depend upon their implicit connections among ideas, memories, emotions, goals, and intentions. Here we briefly the ‘theory of mind’ of users and LLMs, and describe the limitations of our model and propose future research directions.
6.1. A Theory of Users’ Minds for LLMs?
In the context of our revised model of interaction with LLMs, there is no need for LLMs to employ a theory of mind in order to understand the user’s intentions. A theory of mind refers to the ability to explain human behavior by attributing mental states to individuals(Meltzoff, 1999; Wellman, 2014). For example, someone who knows you can create a theory of your mind, including what you know and don’t know about an event, and use it to explain why you answered a question as you did. The theory of mind addresses differences in knowledge and beliefs in individuals when one happens to know how knowledge differs. As discussed above, LLMs are not creating a model of the human when in conversation in even a single session, let alone constructing a model of what a user knows and what they are interested in knowing over time. Even for human experts, it is very challenging – even with knowledge about a person’s past experiences and current knowledge in memory – to model why they may be asking a question now.
Generating a theory of mind sufficient to understand why a person is producing a series of questions (prompts) is a touchstone task for human intelligence. For LLMs, this is beyond the pale. There is no memory generated about the processing the LLM is doing, as all memory is stored through the initial machine learning training and not updated. Additional “fixes” added, such as to avoid the expression of objectionable biases, exist through a separate human-led process of correction entirely apart from the LLM’s knowledge. Most importantly, modeling what humans are thinking is beyond LLMs because the learning algorithms have no goals or intentions of their own. To understand how another mind is understanding, one may require the ability to simulate differences in one’s own thinking; for example, I used to believe this, but now I believe something else. LLMs are not able to recognize contradictions in their own generated outputs, nor to adapt to such experiences. Because these models are probabilistic at their core, they possess no inherent sense of goals, intentions, or feelings. Thus, the fact that they can produce a response in line with humans in a conversation is attributed to their ability to output what is most frequently associated within the training corpus. LLMs are responsive to human intentions in inputs because their data – not their processes – captures the intentions of the humans who created the training corpus. The mathematical solutions tying words to next words implicitly encode the meaning of the original writers who generated those texts, and, therefore the connections between goals and their execution in text output.
6.2. A Theory of LLM Minds for Users?
Does the user need a theory of mind for the LLM? That presumption is the standard approach in HCI, where the user’s mental model provides a functional account of how the system produces its output, and that knowledge is used to generate the necessary action specifications and to evaluate whether the output is ”as good as possible.“ Few humans understand the LLMs well enough to describe what happens at a process level, and that understanding does not allow the prediction of output from input. The reason is that the source of connections between them arises not from the computations performed by the LLM but from the database of texts used in training the model. The LLM itself does not have goals, intentions, or mental states like beliefs, or feelings, but a probabilistic mathematical process that can be explained in a comprehensible way yet will be unsatisfying for understanding why the outputs that it provides to prompts often resemble the kind of conversational turn that a person would produce.
@@ 6.2. A Theory of LLM Minds for Users?
The requisite theory of mind to account for an LLM’s performance includes both the training process and the algorithms used to learn; further, it must include a theory of what was learned by the LLM. For another human, I can use my own mind to generate predictions of what output will likely be produced in another mind. The colors red, white, and blue make me think of America, and that will occur for some other minds as well. However, no other minds have the equivalent dataset of an LLM. Its scope defies the predictions one can generate from the information encoded in one human mind. Without a sense of the outcome of that learning process, it is impossible for another mind to predict its output. That is why LLMs are exciting as generators different from humans but connected through text descriptions of human experiences. It is sufficient to engage in a conversational LLM interaction and feel like talking with another person who has a theory of mind about the individual and associated beliefs, feelings, and goals through the power of intentions.
@@ 6.3. Limitations and Future Work
Of course, LLMs have changed the game for AIs, and they already have a wildly diverging portfolio of application. They make clear the fact that much of human behavior is rote, standardized, and repetitive. Strong patterns across the text database reveal how little originality exists in the accumulated human text products available in digital form. It is quite possible that the deep cognitive processing proposed here as a means of probing LLMs to produce output more similar to desired outcomes is much more rare. However, its value when it does occur suggests it is well worth developing models of the cognitive processes shared across tasks.
@@
Another limitation is the obvious differences among people in their ability to identify intentions from goals. Divergent thinking is a minor part of academic training across the school curriculum where converging on a single correct answer dominates learning. People show major differences in ability to solve open-ended problems and complete generative tasks, potentially attributable to differences in cognitive capacities including memory but also in imagination. A common test of creativity, the unusual uses test, asks for different ways to use a common object such as a brick. People often fixate on functions such as using the brick as a weight, and generate uses like an anchor, paperweight, and balloon holder. Less often, they identify unusual functions, such as using its material as a dye for crayons or lipstick. Perhaps using human minds as the ”key“ limits solutions to only those generated by the human mind. While the LLMs are currently making use of the text-based products of human minds, it is possible that future systems will encode different forms of data not as dependent on goals and intentions. Further, more systems with corpuses combining products of humans and AIs may more further away from cognitive models as explanations of links between input and output.
@@
Finally, this work is intended to propose a direction for the development of new approaches to HCI. Further pathways for Human-LLM interaction can be identified, and new interface supports for LLM use based on prompting guidance are growing daily. While other approaches may be quicker (and dirtier) so as to plug obvious holes in current LLMs, the promise of this work is to capture the intentions found useful by humans in executing generative tasks. To test this approach, comparing prompts where intention is evident will determine its value in creating satisfying outputs from LLMs. Strategies from cowork, such as asking someone to repeat back their understanding of the task instructions given, may prove similarly useful with LLMs. An empirical agenda can determine not just the factual or writing quality of AI systems, but their value to human users. This proposed approach to HCI with LLMs aims to support the user as they must think more deeply and fully during interactions with systems in order to integrate their processing abilities with strengths offered by systems. This work, like development of LLMs, is at its beginning stages.
@@ 7.1. Documented Challenges of LLMs
A core challenge of using LLMs is its explainability (Mathews, 2019), i.e., how can we explain why a model behaves in a certain manner? Compared to the domain of traditional machine learning, LLM explainability is a different challenge (Zhao et al., 2023), as these models are pre-trained to do a variety of complex reasoning tasks (Yang et al., 2023) and absorb patterns from data automatically (Mirchandani et al., 2023). Regarding inputs to LLMs (i.e., prompts), the largest issue is that it is not always clear how a prompt strategy affects model output (Liu et al., 2021b; Sanh et al., 2022). Even for popular methods like chain-of-thought – which asks a model to explain itself – there is no evidence to suggest whether models reason towards the answer through the steps they provide, based on their pre-training data, or through other heuristics (Saparov and He, 2022). Overall, explainability, or the lack thereof, is a significant contributor to the gaps involved in LLM interaction, as users struggle to build a mental model (Bhatt et al., 2021; Sun et al., 2022; Vasconcelos et al., 2023).
@@ 7.2. Prompt Engineering
Another predominant technique used in prompting is chain-of-thought, or providing a series of reasoning steps to show the model how to get to the final answer (Wei et al., 2022; Kojima et al., 2023). These prompts are helpful to learning because for humans, explanations break down why a certain answer is correct and not just what the final answer is (Lampinen et al., 2022; kyoung Ahn et al., 1992; Lombrozo and Carey, 2006; Prystawski et al., 2023). Coupled with few-shot prompting, there are numerous factors that make for useful chain-of-thought prompts, such as the amount of complexity within the prompts (measured by the number of steps), whether the provided examples are relevant to future queries, and the correct ordering of the reasoning steps (Fu et al., 2023; Wang et al., 2023b; Chen et al., 2023b). There are also a plethora of variations that build upon chain-of-thought, including sampling multiple responses given the same prompt (Wang et al., 2023c; Stanovich and West, 2000) (Self-Constency Sampling); repeatedly prompting a language model to ask follow-up questions (Press et al., 2023) (Self-Ask); and both decomposing a problem into numerous steps and sampling numerous responses at each step (Yao et al., 2023) (Tree of Thoughts).
7.3. Designing Natural Language Interfaces
NLI systems enable users to select, control, and change objects within a system using human language. These interfaces are most prevalent in the domains of data visualization (Srinivasan and Stasko, 2020; Gao et al., 2015) and querying (Affolter et al., 2019; Őzcan et al., 2020). There are several issues involved in the development of NLI systems, including the ambiguity of natural language, communicating to users what the system can do, and evaluating the utility of these interfaces in accomplishing their end goal (Shen et al., 2022). To this end, researchers within academia and industry have put forth ideas for creating effective NLI systems, including the use of autocomplete to show users how to phrase their queries to the system (Bacci et al., 2020; Yu and Silva, 2019), the design of conversational interfaces to engage users in a back-and-forth dialogue about their intentions (Setlur and Tory, 2022; Setlur et al., 2016; Hoque et al., 2018), and the development of multi-modal features to give users more control over their input to the interface (Kassel and Rohs, 2018; Srinivasan et al., 2020).
@@ 7.3. Designing Natural Language Interfaces
Going beyond NLI systems, the rise of artificial intelligence (AI) has led to a surge of interest in the development of human-AI systems. The creation of these novel systems has brought new challenges. For instance, since these models are sometimes perceived as non-deterministic “black-boxes,” users can have a hard time discerning how these interfaces produce their outputs (Hoffman et al., 2018; Bathaee, 2017). Likewise, given the massive amount of data these models are trained on, there are new concerns around the bias and privacy of such systems (Bolukbasi et al., 2016; Kostick-Quenet et al., 2022; Yeom et al., 2018). Perhaps most critical is the change in the relationship between humans and interfaces: while traditional NLI systems served more as assistants to end users (Farooq and Grudin, 2016; Wickens et al., 2015), human-AI systems act more as collaborators due to their human-like cognitive abilities (Lyons et al., 2018; Brandt et al., 2018). As a result of this role change, there are many propositions for human-AI interaction guidelines (Amershi et al., 2019; Mohseni et al., 2021; Wright et al., 2020), including conveying the consequences of user actions, providing diverse options from models, and holding the system accountable for errors.
@@
Finally, a lot of recent work in HCI has been focused on generative AI, in particular, LLMs (OpenAI, 2023a; Google, 2023; Anthropic, 2023) and diffusion models (Ramesh et al., 2022; Rombach et al., 2022). Within this space, significant effort has been concentrated on developing effective prompt strategies (see Section  7.2), helping users craft and author these prompts (Zamfirescu-Pereira et al., 2023; Fiannaca et al., 2023; Wu et al., 2022b), and also developing preliminary guidelines for generative models (Chen et al., 2023a; Weisz et al., 2023). Focusing specifically on LLMs, though, what is missing in the current literature is a better understanding of why human-LLM interaction is different from human-AI interaction – and more broadly, how human-LLM systems are different from traditional natural language interfaces – and what types of strategies designers can employ to address the gaps that result from these new challenges.

### v2

@@ 6. Discussion
In this work, we have theorized about cognitive challenges emerging in the transition from conventional software paradigms to prompt-based interactions powered by generative models. Based on prior empirical evidence on challenges with prompting (Zamfirescu-Pereira et al., 2023a, b; Kim et al., 2023), we have applied cognitive science and HCI perspectives to characterize significant HCI design challenges with prompt-based interactions. Given the advanced cognitive capabilities of LLMs, people are now able to express in natural language their bespoke task goals and ask the LLM to perform those goals for them. At the same time, they lack the specific affordances of conventional systems in formulating their intentions and task plans and evaluating the LLM outputs. Given the shift in the operational scope from deterministic functions to dynamic intelligent agents, we have identified new cognitive process models for specifying actions through intentions, i.e., the process of envisioning. In reasoning about envisioning intentions with LLMs, we have also identified three specific gaps including the capability gap, instruction gap, and intentionality gap, and we have provided initial recommendations for interface designers to scaffold prompting. However, a number of open questions remain about designing prompt-based interfaces. Here, we propose open questions for future research as we consider future development and applications of generative models.
6.1. Open Questions for Designing Human-LLM Interactions
6.1.1. How should we model conversational interactions between humans and LLMs?
Given the human-like conversational abilities of LLM interactions, designers must consider how to effectively model conversations. Recent work on designing natural language interfaces (e.g., (Setlur and Tory, 2022; Dombi et al., 2022) ) has primarily considered the ‘Recipient Design’ approach based on Grice’s Cooperative principles (Grice, 1975; Stone, 2005; Branigan et al., 2010). This approach models communication as sensitive to context and individual characteristics, and it recognizes that speakers tailor their speech and communicative behavior to meet the needs of their listeners. As the influential literary theorist Mikhail Bahktin theorizes  (Bakhtin, [n. d.]), any utterance has both addressability (every word always addressed to someone) and answerability (every word directed toward and anticipating an answer from someone). Sociolinguists describe these features as manifest in the recipient design of an utterance  (Goodwin and Heritage, 1990). The maxims of quantity, quality, relation, and manner guide how information is conveyed and interpreted to ensure clarity, relevance, and truthfulness. However, with LLM dialog, such an inductive approach can be challenging for users as these maxims depend on each conversationalist tracking a theory of mind (Frith and Frith, 2005) regarding what their partner thinks and knows. An AI capable of a complex theory of mind for individual users across sessions will likely require extensive development and testing for feasibility. An alternative model is the Transactional Model of Communication (Sameroff, 2009), which offers a more dynamic view of interchange. Rather than precision, a deductive process allows for repairs and adjustments of misunderstanding and miscommunication, and consider the contextual and continuous nature of communication. Applying this model for human-LLM interaction, rather than emphasizing the quality of communication, in interfaces offers opportunities for repair and feedback, e.g., editing a previous prompt in the ChatGPT interface.
Other models of communication, such as the Socio-Cognitive approach to Pragmatics (Kecskes, 2010) and Speech Act Theory (Searle et al., 1980), may also be useful for design. The socio-cognitive approach accounts for how the speaker’s and listener’s background, intentions, and situational contexts contribute to constructing meaning. This approach may provide guidance for characterizing the agentic roles of LLMs in personalization and adaptability, modeling intentions and expectations, code-switching, and ethical and responsible interactions. Further, focusing on the ‘functions’ of language, i.e., speech acts, may allow design for conversational rules and conventions. Lastly, according to social exchange theory (Cropanzano and Mitchell, 2005), communication behaviors are influenced by design to maximize benefits and minimize the cost of interactions. Future research should investigate the balance and trade-offs between high-quality intuitive prompting and the adaptability offered by more deductive conversational approaches. This exploration should focus on how varying models of communication can enhance LLMs’ ability to understand and adapt to diverse user backgrounds, intentions, and contexts. By integrating insights from these communication models, HCI research can develop LLM interfaces that not only facilitate efficient exchanges but also support ethical, personalized, and contextually sensitive interactions.
6.1.2. What is a useful theory of mind for LLMs to support effective envisioning?
@@
In conventional systems, users’ mental models of systems provide a functional account of how the system produces its output, and that knowledge is used to generate necessary action specifications and evaluate whether the output is “ good enough.” In the case of LLMs, novice users likely do not have a system mental model that can sufficiently describe what happens at a process level, or that understanding does not allow the prediction of output from input. The requisite mental model to account for an LLM’s performance includes both the training process and the algorithms used to learn; further, it must include a theory of what was learned by the LLM. If we consider LLMs from a purely human-like perspective, as in human-human communication, we require a cognitive understanding of how communication works, including forming assumptions and expectations, norms, shared knowledge, feedback processing, symbolic understanding, perspective taking, reciprocity and feedback processing, etc. For another human, I can use my own mind to generate predictions of what output will likely be produced by another mind using my own mind as a guide; for example, the colors red, white, and blue make me think of America. However, no other minds have the equivalent dataset of an LLM. Its scope defies the predictions one can generate from the information encoded in one human mind. Without a sense of the outcome of its learning process, it is impossible for another mind to predict LLM output. That is why LLMs are exciting generators different from humans but connected through text descriptions of human experiences. It is sufficient to engage in a conversational LLM interaction and feel like talking with another person who has a theory of mind about the individual and associated beliefs, feelings, and goals through the power of intentions.
Future research should explore how to bridge the gap between the complex, often opaque inner workings of LLMs and the intuitive understanding of its users. This work will involve creating models that accurately represent the LLM’s operations in a user-friendly manner, aligning them with the system’s actual functionality. The challenge lies in simplifying the complex mechanisms of LLMs without sacrificing essential details that users need to predict the outcomes of their interactions. Prior research on end-user programming has studied the challenges novices face in envisioning simple interactive features and ways to support requirement specification, debugging, and verification (Ko et al., 2004). Similarly, for LLMs, theory should account for the evolving nature of user requirements and the emergent design process, acknowledging that users often learn and refine their understanding of LLMs through experience. Efforts should be concentrated on enhancing the visibility of the LLM’s processing pathways, perhaps through interactive visualizations or simplified explanatory frameworks to reinforce user understanding through better prompting, iteration, and evaluation.
6.1.3. What is the optimal “sweet spot” for Human-LLM interaction along the three interaction dimensions?
Reconsidering the three dimensions in Figure 1, LLMs are appealing because the inputs can be abstract, complex, and vague (i.e., underspecified), and they can still produce outputs that are good. As mentioned earlier, if the quality of the answer matters, end users will want to iteratively explore the generative features to reach a high-quality answer. In such cases, envisioning can be challenging, and in both our revised model and as seen in prior work (Zamfirescu-Pereira et al., 2023a), we need better guidance in designing interfaces to make envisioning easier for users. When powering interfaces with LLM capabilities, how do we balance Intent Specificity, Functional Flexibility, and Output Determinacy in ways that leverage the strengths of the LLM while aligning with the user’s needs for quality and relevance? For Intent Specificity, a semi-structured (templated) approach is advantageous. While LLMs excel at interpreting and generating responses to open-ended queries, a certain level of specificity in the user’s intent can guide the LLM to produce more targeted and pertinent content. Functional Flexibility should be oriented towards dynamic responsiveness. The LLM’s ability to pivot across different tasks and domains should be fully utilized within a framework that is shaped by the context of the interaction. Future research should investigate efficient pathways and optimal points along these three dimensions. Such inquiries will be likely to include novel interfaces that suggest modifications to vague inputs and feedback systems that learn from each interaction to refine future responses. Moreover, future research should also consider the evolving nature of user expectations and the continuous advancements in LLM capabilities, ensuring that the “sweet spot” for AI and interface design remains a moving target that adapts to the growing sophistication of both users and technology.
@@ 6.2. Limitations and Future Work
Of course, LLMs have been game-changing for AI, and have launched a wildly diverging portfolio of applications. They evidence the fact that much of everyday human task performance is rote, standardized, and repetitive. Strong patterns across the text database reveal how little originality exists in the accumulated human text products available in digital form. It is quite possible that the deep cognitive processing proposed here as a means of probing LLMs to produce output more similar to desired human outcomes is rare. However, its value when it does occur suggests it is well worth developing models of the cognitive processes shared across cognitive tasks.
@@
Another limitation of our model is the obvious differences among people in their ability to identify intentions from goals. Divergent thinking  (Guilford, 1956) is a minor part of academic training across the school curriculum, whereas converging on a single correct answer dominates learning. People show major differences in their ability to solve open-ended problems and complete generative tasks, potentially attributable to differences in cognitive capacities, including memory but also imagination. A common test of creativity, the unusual uses test, asks for different ways to use a common object, such as a brick (Guilford, 1956). People often fixate on functions such as using the brick as a weight, and generate uses like an anchor, paperweight, and balloon holder. Less often, they identify unusual functions, such as using its material as a dye for crayons or lipstick. Perhaps using human minds as the “key” limits solutions to only those generated by the human mind. While the LLMs are currently making use of the text-based products of human minds, it is possible that future systems will encode different forms of data less dependent on human goals and intentions and more content generated by AI models. Further, more systems with corpora combining products of humans and AIs may diverge further away from cognitive models as explanations of links between input and output.
@@
Finally, this work is intended to propose a direction for the development of new approaches to HCI. Further pathways for human-LLM interaction can be identified, and new interface supports for LLM use based on prompting guidance are growing daily. While other approaches may be quicker (and dirtier) so as to plug obvious holes in current LLMs, the promise of this work is to capture the intentions found useful by humans in executing generative tasks. To test this approach, comparing prompts where the intention is evident will determine its value in creating satisfying outputs from LLMs. Strategies from co-work, such as asking someone to repeat back their understanding of the task instructions given, may prove similarly useful with LLMs. An empirical agenda can determine not just the factual or writing quality of AI systems but also their value to human users. This proposed approach to HCI with LLMs aims to support the user as they must think more deeply and fully during interactions with systems in order to integrate their processing abilities with the strengths offered by systems. This work, like the development of LLMs, is at its beginning stages.
@@ 7.1. Documented Challenges of LLMs
A core challenge of using LLMs is their explainability (Mathews, 2019), i.e., how can we explain why a model behaves in a certain manner? Compared to the domain of traditional machine learning, LLM explainability is a different challenge (Zhao et al., 2023), as these models are pre-trained to do a variety of complex reasoning tasks (Yang et al., 2023) and absorb patterns from data automatically (Mirchandani et al., 2023). Regarding inputs to LLMs (i.e., prompts), the largest issue is that it is not always clear how a prompt strategy affects model output (Liu et al., 2021b; Sanh et al., 2022). Even for popular methods like chain-of-thought – which asks a model to explain itself – there is no evidence to suggest whether models reason towards the answer through the steps they provide, based on their pre-training data, or through other heuristics (Saparov and He, 2022). Overall, explainability, or the lack thereof, is a significant contributor to the gaps involved in LLM interaction, as users struggle to build a mental model (Bhatt et al., 2021; Sun et al., 2022; Vasconcelos et al., 2023).
@@ 7.2. Prompt Engineering
Another predominant technique used in prompting is chain-of-thought, or providing a series of reasoning steps to show the model how to get to the final answer (Wei et al., 2022; Kojima et al., 2023). These prompts are helpful to learning because, for humans, explanations break down why a certain answer is correct and not just what the final answer is (Lampinen et al., 2022; kyoung Ahn et al., 1992; Lombrozo and Carey, 2006; Prystawski et al., 2023). Coupled with few-shot prompting, there are numerous factors that make for useful chain-of-thought prompts, such as the amount of complexity within the prompts (measured by the number of steps), whether the provided examples are relevant to future queries, and the correct ordering of the reasoning steps (Fu et al., 2023; Wang et al., 2023b; Chen et al., 2023b). There are also a plethora of variations that build upon chain-of-thought, including sampling multiple responses given the same prompt (Wang et al., 2023c; Stanovich and West, 2000) (Self-Constency Sampling); repeatedly prompting a language model to ask follow-up questions (Press et al., 2023) (Self-Ask); and both decomposing a problem into numerous steps and sampling numerous responses at each step (Yao et al., 2023) (Tree of Thoughts).
7.3. Designing Human-AI Interfaces
@@
The rise of artificial intelligence (AI) has led to a surge of interest in the development of human-AI systems. The creation of these novel systems has brought new challenges. For instance, since these models are sometimes perceived as non-deterministic “black-boxes,” users can have a hard time discerning how these interfaces produce their outputs (Hoffman et al., 2018; Bathaee, 2017). Likewise, given the massive amount of data these models are trained on, there are new concerns around the bias and privacy of such systems (Bolukbasi et al., 2016; Kostick-Quenet et al., 2022; Yeom et al., 2018). Perhaps most critical is the change in the relationship between humans and interfaces: while traditional NLI systems served more as assistants to end users (Farooq and Grudin, 2016; Wickens et al., 2015), human-AI systems act more as collaborators due to their human-like cognitive abilities (Lyons et al., 2018; Brandt et al., 2018). As a result of this role change, there are many propositions for human-AI interaction guidelines (Amershi et al., 2019; Mohseni et al., 2021; Wright et al., 2020), including conveying the consequences of user actions, providing diverse options from models, and holding the system accountable for errors. Further prior research has looked at changing design practices to accommodate the new challenges of designing user experiences for machine learning capabilities (Yang et al., 2018, 2020; Subramonyam et al., 2021, 2022).
Specific to natural language interfaces, prior work has looked at designing natural language interfaces and LLM-powered chatbots (Yang et al., 2019; Zamfirescu-Pereira et al., 2023a). These interfaces are most prevalent in the domains of data visualization (Srinivasan and Stasko, 2020; Gao et al., 2015) and querying (Affolter et al., 2019; Őzcan et al., 2020). There are several issues involved in the development of NLI systems, including the ambiguity of natural language, communicating to users what the system can do, and evaluating the utility of these interfaces in accomplishing their end goal (Shen et al., 2022). To this end, researchers within academia and industry have put forth ideas for creating effective NLI systems, including the use of autocomplete to show users how to phrase their queries to the system (Bacci et al., 2020; Yu and Silva, 2019), the design of conversational interfaces to engage users in a back-and-forth dialogue about their intentions (Setlur and Tory, 2022; Setlur et al., 2016; Hoque et al., 2018), and the development of multi-modal features to give users more control over their input to the interface (Kassel and Rohs, 2018; Srinivasan et al., 2020).
@@
Finally, recent work in HCI has been focused on generative AI, in particular, LLMs (OpenAI, 2023a; Google, 2023; Anthropic, 2023) and diffusion models (Ramesh et al., 2022; Rombach et al., 2022). Within this space, significant effort has been concentrated on developing effective prompt strategies (see Section  7.2), helping users craft and author these prompts (Zamfirescu-Pereira et al., 2023b; Fiannaca et al., 2023; Wu et al., 2022b), and also developing preliminary guidelines for generative models (Chen et al., 2023a; Weisz et al., 2023). Focusing specifically on LLMs, though, what is missing in the current literature is a better understanding of why human-LLM interaction is different from human-AI interaction – and more broadly, how human-LLM systems are different from traditional natural language interfaces – and what types of strategies designers can employ to address the gaps that result from these new computational opportunities.